OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 + 2x + y2 + 4y = 20 (5 points) Select one: a. The center is located at (1, 2), and the radius is 25. b. The center is located at (-1, 2), and the radius is 25. c. The center is located at (-1, -2), and the radius is 5. d. The center is located at (1, 2), and the radius is 5.

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@robtobey @ganeshie8 @gudden

ganeshie8 (ganeshie8):

complete the square for \(x\) terms complete the square for \(y\) terms

ganeshie8 (ganeshie8):

get the equation into form : \(\large \mathbb{(x-h)^2+(y-k)^2 = r^2}\)

ganeshie8 (ganeshie8):

center = \(\large \mathbb{(h,k)}\) radius = \(\large \mathbb{r}\)

OpenStudy (anonymous):

what is that formula called

ganeshie8 (ganeshie8):

thats the equation for standard form of \(circle\)

OpenStudy (anonymous):

is the radius 5?

ganeshie8 (ganeshie8):

if u want answer : http://www.wolframalpha.com/input/?i=x2+%2B+2x+%2B+y2+%2B+4y+%3D+20+&dataset=

ganeshie8 (ganeshie8):

c

ganeshie8 (ganeshie8):

but im sure u wanto knw HOW to solve as well ? :)

OpenStudy (anonymous):

oh i knew it was C thnx can you help me on some more please

ganeshie8 (ganeshie8):

ok, but i feel u still have problems wid 'completing the square' method ?

OpenStudy (anonymous):

maybe we will see:)

ganeshie8 (ganeshie8):

i mean, did u get, how to complete the square ha ? or simply rushing after answers lol

OpenStudy (anonymous):

yes i did i just need checking and ill tag you for the next question ok

ganeshie8 (ganeshie8):

ok thats good :) in the next question show ur work also, so that i wont be asking u so many questions like this ok ?

OpenStudy (anonymous):

ok